Help with this problem?
Find the values of the constants C and b so that the curve y = Cb^(x) contains the points (2 , 16) and (3 , 32)
b = ?? C = ??
Thank you for any explanation!

Answer by JoelSchwartz(130)

(x=2,y=16) (x=3,y=32)
16=Cb^2
32=Cb^3
2Cb^2=Cb^3
2Cb^2/C=Cb^3/C
2b^2/b^2=b^3/b^2
2=b
16=Cb^2
4C=16
C=4